A particle with a charge of 4.0 A particle with a charge of 4.0 C has a mass of 5 C has a mass of 5.0 x 10-3 kg. What electric field directed upward will exactly balance the weight of the particle?
Answer #1

Electric force should be equal to its weight

Fe = Fg

qE = mg

E = mg/q

q = 4.0x 10-6 C

m = 5.0 x 10-3 kg

g = 9.81 m/s2

E = \frac{\left ( 5.0\times 10^{-3} \right )\left ( 9.81 \right )}{4.0\times 10^{-6}}

E =12.26\times 10^{3} \; N/C